Comparative financial statements for Weller Corporation, a merchandising company, for the year ending December 31 appear below. The company did not issue any new common stock during the year. A total of 600,000 shares of common stock were outstanding. The interest rate on the bond payable was 10%, the income tax rate was 40%, and the dividend per share of common stock was $0.75 last year and $0.40 this year. The market value of the company’s common stock at the end of the year was $28. All of the company’s sales are on account.

Weller Corporation
Comparative Balance Sheet
(dollars in thousands)
This Year Last Year
Assets
Current assets:
Cash $ 1,140 $ 1,410
Accounts receivable, net 9,500 6,700
Inventory 13,500 11,600
Prepaid expenses 660 540
Total current assets 24,800 20,250
Property and equipment:
Land 9,800 9,800
Buildings and equipment, net 48,570 39,980
Total property and equipment 58,370 49,780
Total assets $ 83,170 $ 70,030
Liabilities and Stockholders' Equity
Current liabilities:
Accounts payable $ 20,200 $ 18,600
Accrued liabilities 1,060 890
Notes payable, short term 250 250
Total current liabilities 21,510 19,740
Long-term liabilities:
Bonds payable 9,500 9,500
Total liabilities 31,010 29,240
Stockholders' equity:
Common stock 600 600
Additional paid-in capital 4,000 4,000
Total paid-in capital 4,600 4,600
Retained earnings 47,560 36,190
Total stockholders' equity 52,160 40,790
Total liabilities and stockholders' equity $ 83,170 $ 70,030
Weller Corporation
Comparative Income Statement and Reconciliation
(dollars in thousands)
This Year Last Year
Sales $ 72,000 $ 66,000
Cost of goods sold 34,000 39,000
Gross margin 38,000 27,000
Selling and administrative expenses:
Selling expenses 10,600 10,100
Administrative expenses 6,500 6,700
Total selling and administrative expenses 17,100 16,800
Net operating income 20,900 10,200
Interest expense 950 950
Net income before taxes 19,950 9,250
Income taxes 7,980 3,700
Net income 11,970 5,550
Dividends to common stockholders 600 600
Net income added to retained earnings 11,370 4,950
Beginning retained earnings 36,190 31,240
Ending retained earnings $ 47,560 $ 36,190

Required:

Compute the following financial data and ratios for this year:

1. Working capital. (Enter answer in thousands.)

2. Current ratio. (Round your answer to 2 decimal places.)

3. Acid-test ratio. (Round your answer to 2 decimal places.)

1. Working capital
2. Current ratio
3. Acid-test ratio
Answer #1

Working capital = Current assets - Current Liabilities

= 24,800 - 21,510

= 3290

Current Ratio = Current assets /Current liabilities

= 24,800/21,510

= 1.15

Acid test ratio = Quick assets/Current liabilities

= (24,800-13,500-660)/21,510

= 0.49